OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 18. What is the value of y when x = 4?

OpenStudy (amistre64):

to vary directly means to make the same move. you reflection in a mirror varies directly with you. in other words, as x increases, y increases; as x gets small, y gets small

OpenStudy (amistre64):

y = k*x^2; where k is called a constant of variation that makes both sides acutally equal

OpenStudy (amistre64):

when x = 3 and y = 18 we get: 18 = k * 3^2 18 = k * 9 18/9 = k k = 2 ; since we know "k" we can use this to find any relation between y and x now

OpenStudy (amistre64):

y = 2 * x^2 when x=4 we get: y = 2 * 4^2 = 2 * 16 = 32

OpenStudy (anonymous):

thank you :)

OpenStudy (amistre64):

youre welcome :)